OpenStudy (anonymous):

Solve the following system of equations. 3x - 2y = 6 6x - 4y = 12 (0, 0) (6, 12) Infinitely Many Solutions No Solutions

OpenStudy (anonymous):

See, if you multiply 2 with first equation, then you will get second equation or not??

OpenStudy (anonymous):

@Nurali No Solution??

OpenStudy (anonymous):

When we get the same equation for both the lines, then they are overlapping lines.. So they must have Infinitely Many Solutions... this is what I think..

OpenStudy (nurali):

3x - 2y = 6---------(1) 6x - 4y = 12--------(2) multiply equation 1 by '2' and subtract 1 and 2. 2[3x-2y=6] 6x-4y=12 6x-4y=12 6x-4y=12 0=0 This is a special case. Notice that if you multiply the first equation by 2 on both side, you get the second equation. Both equations are a line, with the same slope and intercept. The two lines are on top of each other. So there are an infinite number of solutions.

OpenStudy (anonymous):

OH so its {0,0} right?

OpenStudy (nurali):

no Infinitely Many Solutions